A body of mass 40kg is given an acceleration of 10ms on a horizontal ground for which coefficient of static friction =0.5. Calculate the force required to accelerate the body(take g=10ms)


1
Expert's answer
2021-02-15T00:55:15-0500

Applying the Newton's Second Law of Motion, we get:


"F-F_{fr}=ma,""F-\mu_smg=ma,""F=m(a+\mu_sg),""F=40\ kg\cdot(10\ \dfrac{m}{s^2}+0.5\cdot10\ \dfrac{m}{s^2})=600\ N."